# Bayesian Multiobject Tracking With Neural-Enhanced Motion and Measurement Models

### v2

@@ V Update Step with Neural-Enhanced Measurement Model
In what follows, we develop the neural enhanced measurement model and the corresponding update step. The neural-enhanced measurement model involves two learnable factors referred to as “affinity factors” and “false positive rejection factors” that refine the conventional MOT measurement model. These factors are computed based on neural networks that use measurement and object-oriented features as input.
@@ V-A Neural-Enhanced Measurement Model
$f_{\mathrm{ne}}({\bm{\mathbf{z}}}^{j}_{k}|\underline{\bm{\mathbf{x}}}^{i}_{k})=\mathpzc{f}^{i,j}_{\mathrm{af},k}\hskip 0.85358ptf({\bm{\mathbf{z}}}^{j}_{k}|\underline{\bm{\mathbf{x}}}^{i}_{k}).\vspace{.5mm}$ (35)
In this formulation, the learnable affinity factor $\mathpzc{f}_{\mathrm{af},k}^{i,j}>0$ is introduced to provide measurement-related information that is statistically independent to the model-based likelihood function $f({\bm{\mathbf{z}}}^{j}_{k}|\underline{\bm{\mathbf{x}}}^{i}_{k})$. Specifically, $\mathpzc{f}_{\mathrm{af},k}^{i,j}$ can capture more complex feature information from measurements that is not provided by the detector or has been idenfitied as too challenging to model statistically. It can be interpreted probabilistically as a surrogate for a statistical model describing the dependence of shape and appearance measurements on the kinematic PO state.
@@
[…]
as well as $q_{\mathrm{ne}}(\underline{\bm{\mathbf{x}}}_{k}^{i},0,a_{k}^{i};\bm{\mathbf{z}}_{k})=q(\underline{\bm{\mathbf{x}}}_{k}^{i},0,a_{k}^{i};\bm{\mathbf{z}}_{k})$ and $v_{\mathrm{ne}}(\overline{\bm{\mathbf{x}}}_{k}^{i},$ $0,b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})=v(\overline{\bm{\mathbf{x}}}_{k}^{i},0,b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})$. The neural enhanced likelihood ratios will replace their corresponding model-based counterpart in (15) and (20). This results in neural-enhanced functions $l_{\mathrm{ne}}(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})$ and $n_{\mathrm{ne}}(\bm{\mathbf{y}}^{j}_{k};\bm{\mathbf{z}}_{k})$ used in the proposed measurement update and the new object initialization step, respectively.
@@ V-B Feature Extraction and Computation of Learnable Factors
The extraction of object and measurement-oriented features is discussed next. At time $k$ and for each legacy PO $i$, an object-oriented kinematic feature $\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{ki},k}\in\mathbb{R}^{d_{\mathrm{ki}}}$ is obtained as the mean of the predicted PDF of PO state $\underline{\bm{\mathbf{x}}}^{i}_{k}$, i.e.,
$\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{ki},k}=\int\underline{\bm{\mathbf{x}}}^{i}_{k}f(\underline{\bm{\mathbf{x}}}^{i}_{k}|\underline{r}^{i}_{k}=1,\bm{\mathbf{z}}_{1:k-1})\hskip 0.85358pt\mathrm{d}\underline{\bm{\mathbf{x}}}^{i}_{k}.$ (39)
@@
To obtain an object-oriented shape feature vector, $\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{sa},k}$, for legacy PO $i$ at time $k$, multiple processing stages are performed. Fig. 4 shows these processing stages for one feature map. First, we use the kinematic information (e.g., position and rotation) from time $k-1$ as well as size information to identify a region of interest (ROI) on each feature map $l\in\{1,\dots,L\}$. Size information is either part of the kinematic state or known a priori. A ROI covers a region around the location of PO $i$ on each feature map corresponding to measurements at the last time step $k-1$. The entries of the $l$th feature map corresponding to the ROI used for further processing are denoted as ¯$\underline{\mathcal{R}}_{l,k-1}^{i}$. A shape feature vector, $\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{sa},k}$, is finally obtained by applying a neural network $\mathrm{NN}(\cdot)$ to all ¯$\underline{\mathcal{R}}_{l,k-1}^{i}$, $l\in\{1,\dots,L\}$, i.e.,
$\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{sa},k}=\big{[}\mathrm{NN}(\underline{\mathcal{R}}_{1,k-1}^{i})^{\mathrm{T}}\hskip 2.84526pt\mathrm{NN}(\underline{\mathcal{R}}_{2,k-1}^{i})^{\mathrm{T}}\cdots\mathrm{NN}(\underline{\mathcal{R}}_{L,k-1}^{i})^{\mathrm{T}}\big{]}^{\mathrm{T}}.$
@@
Measurement-oriented features are obtained similarly. In particular, for measurement $j$, a measurement-oriented kinematic feature, $\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{ki},k}$, is directly provided by the output of the detector at time $k$. In addition, a measurement-oriented shape feature $\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{sa},k}$ is extracted by (i) using feature maps corresponding to measurements from discrete time step $k$, (ii) following the same steps as for object-oriented shape features, and (iii) using the kinematic feature $\overline{\bm{\mathbf{f}}}^{i}_{\mathrm{ki},k}$ to identify regions of interest, ¯$\overline{\mathcal{R}}_{l,k}^{j}$, on the feature maps $l\in\{1,\dots,L\}$. Typically, neural networks used to extract object and measurement-oriented shape features share the same weights, ensuring consistency across both processes.
TABLE I: Performance on Nuscenes Test Set.

### v3

@@ V Update Step with Neural-Enhanced Measurement Model
In what follows, we develop the neural enhanced measurement model and the corresponding update step. The neural-enhanced measurement model involves two learnable factors referred to as “affinity factors” and “false positive rejection factors” that refine the conventional MOT measurement model. These factors are computed based on neural networks that use measurement and object-oriented features as
input.
@@ V-A Neural-Enhanced Measurement Model
$f_{\mathrm{ne}}({\bm{\mathbf{z}}}^{j}_{k}|\underline{\bm{\mathbf{x}}}^{i}_{k})=\mathpzc{f}^{i,j}_{\mathrm{af},k}\hskip 0.85358ptf({\bm{\mathbf{z}}}^{j}_{k}|\underline{\bm{\mathbf{x}}}^{i}_{k}).\vskip 1.42262pt$ (26)
In this formulation, the learnable affinity factor $\mathpzc{f}_{\mathrm{af},k}^{i,j}>0$ is introduced to provide measurement-related information that is statistically independent to the model-based likelihood function $f({\bm{\mathbf{z}}}^{j}_{k}|\underline{\bm{\mathbf{x}}}^{i}_{k})$. Specifically, $\mathpzc{f}_{\mathrm{af},k}^{i,j}$ can capture more complex feature information from measurements that is not provided by the detector or has been identified as too challenging to model statistically. It can be interpreted probabilistically as a surrogate for a statistical model describing the dependence of shape and appearance measurements on the kinematic PO state.
@@
$\hskip-5.69054ptq_{\mathrm{ne}}(\underline{\bm{\mathbf{x}}}_{k}^{i},1,a_{k}^{i}=j;\bm{\mathbf{z}}_{k})=$ $~\mathpzc{f}^{j}_{\mathrm{fpr},k}\hskip 0.85358pt\hskip 0.85358pt\hskip 0.85358pt\mathpzc{f}^{i,j}_{\mathrm{af},k}\hskip 0.85358pt\hskip 0.85358ptq(\underline{\bm{\mathbf{x}}}_{k}^{i},1,a_{k}^{i}=j;\bm{\mathbf{z}}_{k})$
$v_{\mathrm{ne}}(\overline{\bm{\mathbf{x}}}_{k}^{j},1,b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})=$ $~\mathpzc{f}^{j}_{\mathrm{fpr},k}\hskip 0.85358pt\hskip 0.85358ptv(\overline{\bm{\mathbf{x}}}_{k}^{j},1,b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})$
as well as $q_{\mathrm{ne}}(\underline{\bm{\mathbf{x}}}_{k}^{i},0,a_{k}^{i};\bm{\mathbf{z}}_{k})\hskip-0.85358pt=\hskip-0.85358ptq(\underline{\bm{\mathbf{x}}}_{k}^{i},0,a_{k}^{i};\bm{\mathbf{z}}_{k})$ and $v_{\mathrm{ne}}(\overline{\bm{\mathbf{x}}}_{k}^{i},$ $0,b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})\hskip-0.85358pt=\hskip-0.85358ptv(\overline{\bm{\mathbf{x}}}_{k}^{i},0,b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})$. The neural enhanced likelihood ratios will replace their corresponding model-based counterpart in (14) and (19). This results in neural-enhanced functions $l_{\mathrm{ne}}(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})$ and $n_{\mathrm{ne}}(\bm{\mathbf{y}}^{j}_{k};\bm{\mathbf{z}}_{k})$ used in the proposed measurement update and the new object initialization step, respectively.
@@ V-B Feature Extraction and Computation of Learnable Factors
The extraction of object and measurement-oriented features is discussed next. At time $k$ and for each legacy PO $i$, an object-oriented kinematic feature $\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{ki},k}\hskip-0.85358pt\in\hskip-0.85358pt\mathbb{R}^{d_{\mathrm{ki}}}$ is obtained as the mean of the predicted PDF of PO state $\underline{\bm{\mathbf{x}}}^{i}_{k}$, i.e.,
$\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{ki},k}=\int\underline{\bm{\mathbf{x}}}^{i}_{k}f(\underline{\bm{\mathbf{x}}}^{i}_{k}|\underline{r}^{i}_{k}=1,\bm{\mathbf{z}}_{1:k-1})\hskip 0.85358pt\mathrm{d}\underline{\bm{\mathbf{x}}}^{i}_{k}.$
@@
To obtain an object-oriented shape feature vector, $\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{sa},k}$, for legacy PO $i$ at time $k$, multiple processing stages are performed. Fig. 4 shows these processing stages for one feature map. First, we use the kinematic information (e.g., position and rotation) from time $k\hskip-0.85358pt-\hskip-0.85358pt1$ as well as size information to identify a region of interest (ROI) on each feature map $l\in\{1,\dots,L\}$. Size information is either part of the kinematic state or known a priori. A ROI covers a region around the location of PO $i$ on each feature map corresponding to measurements at the last time step $k\hskip-0.85358pt-\hskip-0.85358pt1$. The entries of the $l$th feature map corresponding to the ROI used for further processing are denoted as $\underline{\mathcal{R}}_{l,k-1}^{i}$. A shape feature vector, $\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{sa},k}$, is finally obtained by applying a neural network $\mathrm{NN}(\cdot)$ to all $\underline{\mathcal{R}}_{l,k-1}^{i}$, $l\hskip-0.85358pt\in\hskip-0.85358pt\{1,\dots,L\}$, i.e.
$\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{sa},k}=\big[\mathrm{NN}(\underline{\mathcal{R}}_{1,k-1}^{i})^{\mathrm{T}}\hskip 2.84526pt\mathrm{NN}(\underline{\mathcal{R}}_{2,k-1}^{i})^{\mathrm{T}}\cdots\mathrm{NN}(\underline{\mathcal{R}}_{L,k-1}^{i})^{\mathrm{T}}\big]^{\mathrm{T}}\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt.$
@@
Measurement-oriented features are obtained similarly. In particular, for measurement $j$, a measurement-oriented kinematic feature, $\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{ki},k}$, is directly provided by the output of the detector at time $k$. In addition, a measurement-oriented shape feature $\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{sa},k}$ is extracted by (i) using feature maps corresponding to measurements from discrete time step $k$, (ii) following the same steps as for object-oriented shape features, and (iii) using the kinematic feature $\overline{\bm{\mathbf{f}}}^{i}_{\mathrm{ki},k}$ to identify regions of interest, $\overline{\mathcal{R}}_{l,k}^{j}$, on the feature maps $l\in\{1,\dots,L\}$. Typically, neural networks used to extract object and measurement-oriented shape features share the same weights, ensuring consistency across both processes.
By using kinematic features and shape features for each PO and measurement pair, $(i,j)$, we can compute the affinity factors $\mathpzc{f}^{i,j}_{\mathrm{af},k}$ for the proposed neural-enhanced measurement model. Specifically, for each $(i,j)$, the affinity factor $\mathpzc{f}^{i,j}_{\mathrm{af},k}$ is obtained
as
$\mathpzc{f}^{i,j}_{\mathrm{af},k}=\mathrm{exp}\big({\mathrm{NN}}\big(\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{ki},k},\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{sa},k},\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{ki},k},\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{sa},k}\big)\big)$
where $\mathrm{NN}(\cdot)$ is a neural network. The exponential mapping guarantees positivity and naturally controls the factor’s dynamic range compared to the corresponding model-based likelihood function $f(\bm{\mathbf{z}}^{j}_{k}|\underline{\bm{\mathbf{x}}}_{k}^{i})$. An optional rectified linear unit (ReLU) could be applied to $\mathpzc{f}^{i,j}_{\mathrm{af},k}$. This unit enforces that $\mathpzc{f}^{i,j}_{\mathrm{af},k}\hskip-0.85358pt\geqslant\hskip-0.85358pt1$, i.e., the affinity factor can only increase the likelihood that measurement $j$ is associated with PO $i$ but not decrease it. It has been observed that applying the ReLU to $\mathpzc{f}^{i,j}_{\mathrm{af},k}$ can significantly reduce the training variance but potentially lead to a slightly reduced overall MOT performance. The false positive rejection factor $f^{j}_{\mathrm{fpr},k}\in(0,1)$ for each measurement $j$, is computed by
$\mathpzc{f}^{j}_{\mathrm{fpr},k}=$ $\sigma\Big[{\mathrm{NN}}\big(\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{ki},k},\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{sa},k}\big)\Big]$
where $\sigma[\cdot]$ is the sigmoid function that restricts the output to $(0,1)$.
VI Numerical Analysis – Scenario and Implementation Details
In this section, we will introduce the autonomous driving scenario used for performance evaluation and discuss the loss function as well as the feature extraction mechansim. A block diagram of the proposed Bayesian MOT framework is shown in Fig. 5
VI-A The NuScenes Autonomous Driving Dataset
In particular, we employ LiDAR and camera data from the nuScenes autonomous driving dataset [10]. This dataset consists of 1000 autonomous driving scenes and seven object classes. Tracking is performed in 2-D. The kinematic state of objects is defined as $\bm{\mathbf{x}}^{i}_{k}=\big[(\bm{\mathbf{p}}^{i}_{k})^{\mathrm{T}}\hskip 0.85358pt(\bm{\mathbf{v}}^{i}_{k})^{\mathrm{T}}\big]^{\mathrm{T}}\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt$, where $\bm{\mathbf{p}}^{i}_{k}\in\mathbb{R}^{2}$ and $\bm{\mathbf{v}}^{i}_{k}\in\mathbb{R}^{2}$ are the 2-D position and velocity, respectively. At each time step $k$, measurements are provided by a detector that consists of a neural network [11, 69]. In particular, for each detected object, measurements of the 2-D position, the 2-D velocity, a reliability score, as well as the size and orientation of the bounding box are provided. The 2-D position and 2-D velocity measurement, denoted as $\bm{\mathbf{z}}^{j}_{k}\in\mathbb{R}^{4}$ for each $j$, is directly used in the statistical model for MOT. In particular, the likelihood function $f(\bm{\mathbf{z}}^{j}_{k}|\bm{\mathbf{x}}_{k}^{i})$ used in (17) and (12) is based on the linear model $\bm{\mathbf{z}}^{j}_{k}=\bm{\mathbf{x}}_{k}^{i}+\bm{\mathbf{r}}_{k}^{j}$ where $\bm{\mathbf{r}}_{k}^{j}\sim\mathcal{N}(\bm{\mathbf{0}};\bm{\mathbf{\Sigma}}_{r})$ and $\bm{\mathbf{\Sigma}}_{r}=\mathrm{diag}\{[\sigma^{2}_{p}\hskip 0.85358pt\hskip 0.85358pt\sigma^{2}_{p}\hskip 0.85358pt\hskip 0.85358pt\sigma^{2}_{v}\hskip 0.85358pt\hskip 0.85358pt\sigma^{2}_{v}]^{\mathrm{T}}\}$. Here, position and velocity-related variances $\sigma^{2}_{p}$ and $\sigma^{2}_{v}$ are determined as discussed later in this section. Size and orientation of the bounding box are used for feature extraction. The reliability score is used for final performance evaluation as also discussed later in this section.
Figure 5: Block diagram of the proposed Bayesian MOT method for a single time step $k$. The PDF computation in the proposed method is based on the indicated factor graph and corresponding message-passing procedure [23, 38]. Different parts of the proposed method are indicated by different colors (1) green: input and output PDF related to time step $k$; (2) blue: computations related to the prediction step (3) red: computations related to the evaluation of the likelihood function (4) black/gray: probabilistic data association. The following shorthand notations are used: $f^{i}=f(\underline{\bm{\mathbf{y}}}^{i}_{k}|{\bm{\mathbf{y}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},{\mathbf{h}}^{i}_{k-1})$, $\hat{q}^{i}=\hat{q}(\underline{\bm{\mathbf{y}}}_{k}^{i},a_{k}^{i};\bm{\mathbf{z}}_{k})$, $\hat{v}^{j}=\hat{v}(\overline{\bm{\mathbf{y}}}_{k}^{j},b_{k}^{j};\bm{\mathbf{z}}_{k})$, $\phi^{i,j}=\phi^{i,j}(a_{k}^{i},b_{k}^{j})$.
Measurements are pre-processed using the non-maximum suppression (NMS) technique [40]. Here, the threshold is set to 0.1, and no score filtering is performed. As discussed in Section IV-C, the prediction step is implemented using SPs and results in a Gaussian representation of each PO state. The update step, on the other hand, is implemented using $10^{4}$ particles to represent each PO state. The Gaussian PDF resulting from the prediction step is used as a proposal PDF for the particle-based update step. From the particles resulting from the update step, a Gaussian representation is extracted by computing the sample mean and sample covariance. This Gaussian representation is then used as an input for the prediction step at the next time step. Using a particle-based implementation despite the linear sensor measurement model has been demonstrated to lead to superior data association results and thus an improved overall tracking performance [22].
Together with the probability of existence, the reliability score is used to compute the final AMOTA for each PO. The detection region of the sensor is given by $p^{x}_{k}-54,p^{x}_{k}+54]\times[p^{y}_{k}-54,p^{y}_{k}+54]$, where $p^{x}_{k},p^{y}_{k}$ represent the time-dependent 2D positions of the ego vehicle. The prior PDF of false positives $f_{\mathrm{fp}}(\cdot)$ and newly detected objects $f_{\mathrm{n}}(\cdot)$ are uniformly distributed over the tracking region.The threshold for object declaration is $T_{\mathrm{dec}}=0.5$ and the threshold for pruning is $T_{\mathrm{pru}}=10^{-4}$ (cf. Section III-C). The survival probability is set to $p_{s}=0.999$. The parameters $p_{d}$, $\sigma^{2}_{p}$, $\sigma^{2}_{v}$, $Q$, $C_{\mathrm{n}}(\cdot)$, $C_{\mathrm{fp}}(\cdot)$ are estimated from the training data. In the neural-enhanced prediction step, we consider a maximum number of $M=10$ neighbors based on their distance to the object to be predicted.
VI-B Loss Function and Training Procedure
The computation of the loss $\ell_{k}$ at time $k$ used for the training of the proposed neural architecture can split up in a component related to the motion model, i.e., $\ell^{\hskip 0.85358pt\mathrm{motion}}_{k}\hskip-0.85358pt$, and a component related to the measurement model, $\ell_{k}^{\mathrm{meas}}$.
The motion-model related component of the loss is given by
[…]
where we recall that $\hat{\bm{\mathbf{x}}}_{k}^{i}$ is the MMSE estimate of the kinematic state of object $i$ at time $k$ and $\bm{\mathbf{x}}^{i}_{\mathrm{gt},k}$ is the corresponding ground truth. For the evaluation of (28), we need to decide which estimated object state can be associated with a ground truth state and only compute the loss for them. This decision is performed by applying the Hungarian algorithm [25]. Since only some kinematic object states are associated with ground truth states, we have that $I^{\prime}_{k}\leq I_{k}$.
For the measurement model-related component of the loss, we use the “weighted” binary cross-entropy loss [5, Chapter 4.3]. In particular, the loss $\ell_{k}^{\mathrm{af}}$ for the learning of affinity factors is given by
$\ell_{1,k}^{\mathrm{af}}=$ $~-\frac{\sum_{i=1}^{I_{k}}\sum_{j=1}^{J_{k}}\mathpzc{f}_{\mathrm{gt},k}^{i,j}\ln\big(\sigma(\ln(\mathpzc{f}_{\mathrm{af},k}^{i,j})))}{\sum_{i=1}^{I_{k}}\sum_{j=1}^{J_{k}}\mathpzc{f}_{\mathrm{gt},k}^{i,j}}$ (29)
$\ell_{2,k}^{\mathrm{af}}=$ $~-\frac{\sum_{i=1}^{I_{k}}\sum_{j=1}^{J_{k}}(1-\mathpzc{f}_{\mathrm{gt},k}^{i,j})\ln\big(1-\sigma(\ln(\mathpzc{f}_{\mathrm{af},k}^{i,j})))}{\sum_{i=1}^{I_{k}}\sum_{j=1}^{J_{k}}(1-\mathpzc{f}_{\mathrm{gt},k}^{i,j})}$ (30)
$\ell_{k}^{\mathrm{af}}=$ $~\ell_{1,k}^{\mathrm{af}}+\ell_{2,k}^{\mathrm{af}}$
where $\sigma(\cdot)$ denotes the sigmoid function and $\mathpzc{f}_{\mathrm{gt},k}^{i,j}\in\{0,1\}$ denotes the ground truth association of measurement $j$ to PO $i$, i.e., if PO $i$ is indeed matched with measurement $j$, we have $\mathpzc{f}_{\mathrm{gt},k}^{i,j}\hskip-0.85358pt=\hskip-0.85358pt1$, otherwise $\mathpzc{f}^{i,j}_{\mathrm{gt},k}\hskip-0.85358pt=\hskip-0.85358pt0$. The loss $\ell_{1,k}^{\mathrm{af}}$ characterizes accuracy of existing associations, while $\ell_{2,k}^{\mathrm{af}}$ characterizes accuracy of missing associations. The normalization terms in (29) and (30) address the data imbalance problem, i.e., the normalization makes sure that whether a PO $i$ is associated with a measurement $j$ or not equally contributes to the joint loss $\ell_{k}^{\mathrm{af}}$, regardless of the specific number of POs and measurements.
For false positive rejection coefficients $\mathpzc{f}_{\mathrm{fpr},k}^{i,j}$, the corresponding loss function is given
by
$\ell_{1,k}^{\mathrm{fpr}}$ $=\frac{\sum_{j=1}^{J_{k}}\mathpzc{f}^{j}_{\mathrm{gt},k}\ln(\mathpzc{f}^{j}_{\mathrm{fpr},k})}{\sum_{j=1}^{J_{k}}\mathpzc{f}^{j}_{\mathrm{gt},k}}$
$\ell_{2,k}^{\mathrm{fpr}}$ $=-w_{\mathrm{\mathrm{fpr}}}\hskip 0.85358pt\hskip 0.85358pt\frac{\sum_{j=1}^{J_{k}}(1-\mathpzc{f}^{j}_{\mathrm{gt},k})\ln(1-\mathpzc{f}^{j}_{\mathrm{fpr},k})}{\sum_{j=1}^{J_{k}}(1-\mathpzc{f}^{j}_{\mathrm{gt},k})}$
$\ell_{k}^{\mathrm{fpr}}$ $=\ell_{1,k}^{\mathrm{fpr}}+\ell_{2,k}^{\mathrm{fpr}}$
where $\mathpzc{f}^{j}_{\mathrm{gt},k}\in\{0,1\}$ represents the ground truth label indicating for each measurement if it is a false positive or not, and $w_{\mathrm{\mathrm{fpr}}}\geqslant 0$ is a parameter that makes it possible to prioritize the loss contribution $\ell_{1,k}^{\mathrm{fpr}}$ over $\ell_{2,k}^{\mathrm{fpr}}$ and vice versa. The loss contribution $\ell_{1,k}^{\mathrm{fpr}}$ represents the classification accuracy for true measurements and $\ell_{2,k}^{\mathrm{fpr}}$ represents the classification accuracy for false positives. The tuning parameter is typically $w_{\mathrm{\mathrm{fpr}}}\hskip-0.85358pt<\hskip-0.85358pt1$. This is motivated by the fact that missing an object is usually more harmful than generating a false positive. Finally, the measurement model-related component of the loss is obtained as $\ell^{\hskip 0.85358pt\mathrm{meas.}}_{k}=\ell_{k}^{\mathrm{af}}+\ell_{k}^{\mathrm{fpr}}$.
Finally, the joint loss is given by
$\ell=\ell^{\hskip 0.85358pt\mathrm{motion}}_{k}+w^{\hskip 0.85358pt\mathrm{meas.}}_{\mathrm{\ell}}\hskip 0.85358pt\ell^{\hskip 0.85358pt\mathrm{meas.}}_{k}$
where $w_{\mathrm{\mathrm{fpr}}}\geqslant 0$ is another parameter that makes it possible to prioritize the loss contribution $\ell^{\hskip 0.85358pt\mathrm{meas.}}_{k}$ over $\ell^{\hskip 0.85358pt\mathrm{motion}}_{k}$ and vice versa. This joint loss function makes it possible to train the neural enhanced motion and measurement models jointly.
The input data for the pretraining of the motion model is generated from ground truth information. In particular, to obtain the input data, random noise is added, deterministic biases are applied, and certain states are randomly removed from tracks. For the training of the measurement model, we add additional Gaussian noise to the measurements. This is motivated by the fact that the measurements generated by detectors in the training dataset are typically of higher quality than those in the validation dataset. The joint training of the proposed method is performed based on the Adam optimizer. The batch size and learning rate are set to 1 and $10^{-4}$. The tuning parameter $w_{\mathrm{\mathrm{fpr}}}$ is 0.1. The pre-training for the transition function is based on the Adam optimizer. The batch size and learning rate are set to 32 and $10^{-3}$. The following strategy has been identified as a means to make the training more stable. First, the motion model is pretrained individually, and then its training is refined by joint training with the measurement model. For the pretraining of the motion model, we can directly use ground-truth tracks and avoid the aforementioned assignment by means of the Hungarian algorithm. We also explored more complex architectures — including deeper/wider MLPs, CNN-based models, and other variants — but observed only marginal differences in performance. Thus, we adopt a rather simple and efficient network structure with an overall depth of 10 layers.
VI-C Feature Extraction
The object-oriented kinematic feature is given by $\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{ki},k}\hskip-0.85358pt=[\underline{\bm{\mathbf{p}}}_{k}^{i\hskip 0.85358pt\mathrm{T}}\hskip 0.85358pt{\underline{\bm{\mathbf{v}}}}^{i\hskip 0.85358pt\mathrm{T}}_{k}\hskip 0.85358pt{\underline{\bm{\mathbf{u}}}}^{i\hskip 0.85358pt\mathrm{T}}_{k}]^{\mathrm{T}}\hskip-0.85358pt\hskip-0.85358pt.$ Here, $\underline{\bm{\mathbf{p}}}_{k}^{i}$ and ${\underline{\bm{\mathbf{v}}}}^{i}_{k}$ are the 2-D position and velocity estimate extracted as the mean of the predicted posterior PDF $f({\bm{\mathbf{x}}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k-1},r^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt1)$ and ${\underline{\bm{\mathbf{u}}}}^{i}_{k}$ represents size and orientation information. In particular, we use bounding box information of the measurement used to initialize PO $i$ as size information. In addition, orientation information is extracted from the last measurement associated with PO $i$. The measurement-oriented kinematic feature $\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{ki},k}\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358pt\hskip-0.85358pt[\overline{\bm{\mathbf{p}}}_{k}^{j\hskip 0.85358pt\mathrm{T}}\hskip 0.85358pt\hskip 0.85358pt{\overline{\bm{\mathbf{v}}}}^{j\hskip 0.85358pt\mathrm{T}}_{k}\hskip 0.85358pt\hskip 0.85358pt{\overline{\bm{\mathbf{u}}}}^{j\hskip 0.85358pt\mathrm{T}}_{k}]^{\mathrm{T}}$ consists of a 2-D position and velocity measurement as well as size and feature information provided by the detector.
TABLE I: Performance Evaluation Based on the Nuscenes Test Set